Karvel Corporation uses a predetermined overhead rate based on machine-hours to apply
manufacturing overhead to jobs. For the month of August, Karvel estimated total manufacturing
overhead costs at £300,000 and total machine-hours at 75,000 hours. Actual results for the period
were manufacturing overhead costs of £290,000 and 75,000 machine-hours. As a result, Karvel
would have:
1. applied more overhead to Work in Process than the actual amount of overhead cost for
the year.
2. applied an amount of overhead to Work in Process that was equal to the actual amount of
overhead.
3. applied less overhead to Work in Process than the actual amount of overhead cost for the
year.
4. found it necessary to recalculate the predetermined overhead rate.
Answer #1

Predetermined Overhead Rate = Total Estimated Manufacturing Overhead / Total Estimated Machine Hours

Total Estimated Manufacturing Overhead = 300,000

Total Estimated Machine Hours = 75,000 Hours

Predetermined Overhead Rate = 300,000 / 75,000

Predetermined Overhead Rate = 4 per Machine Hour

Manufacturing Overhead Applied = Predetermined Overhead Rate * Actual Machine Hours Worked

Manufacturing Overhead Applied = 4 * 75,000

Manufacturing Overhead Applied = 300,000

Manufacturing Overhead Overapplied (Underapplied) = Total Overhead Applied - Actual Overhead

Manufacturing Overhead Overapplied (Underapplied) = 300,000 - 290,000

Manufacturing Overhead Overapplied = 10,000

Manufacturing Overhead are Overapplied by 10,000.

Option 1 is the Correct Answer.

Karvel would have applied more Overhead to Work in Process than the actual amount of Overhead Cost for the year.